Calculate the value of the thermodynamic equilibrium (K) constant at 358 K

Cl2(g) + Br2(l) ⟶ 2 BrCl(g)
ΔfH° at 298.15 K

/ kJ mol−1

0 0 14.68
Sm° at 298.15 K

/ J K−1 mol−1

223.1 152.2 240.1
1
Expert's answer
2018-03-10T04:02:07-0500
∆H°(rxn)= 2*14.68-(0+0)=29.36 kJ/mol
∆S°(rxn)=2*240.1-(223.1+152.2)=104.9 J/K•mol
∆G°=∆H°-T∆S°
∆G°=29360J - (358K)*(104.9J/K) = - 8194.2 J
K = exp(-∆G°/RT) = exp(-(-8194.2J)/8.314•(358K)) = exp(2.753) = 15.69